Can fermats principle be used to derive Rayleigh Criteria of Resolution
1
Expert's answer
2020-04-15T10:26:12-0400

Fermat Principle:- A ray of light transverse from on point to another point by a rout which takes least time.

momentum of photon "(p)=\dfrac{h\nu}{c}"

Rayleigh Criteria: The Rayleigh criterion stated in the equation "\u03b8=1.22\u03bbD" gives the smallest possible angle θ between point sources.

The intensity of the expression for grating for peak at the phase "(\delta)=2m\pi"

The closest minimum "(\delta)=\pi/2"

Change in phase "\triangle \delta=\dfrac{2\pi}{\lambda}b\sin\theta"

Now, taking the partial differentiation,

"d\delta=\dfrac{2\pi}{\lambda}b\cos \theta d\theta"

For maximum condition,

"b\sin\theta = m\lambda"

differential equation,

"b\cos\theta d\theta=md\lambda"


now substituting the value in the above equation,

"\dfrac{2\pi}{N}=\dfrac{2\pi}{\lambda}md\lambda"

Now, by rearranging,

"R=\dfrac{\lambda}{\Delta\lambda}=mN"